Write the value of following 1) vector j .vector k 2)vector k . Vector k

Grade:11

1 Answers

Dhananjay
6 years ago
Angle between vector k and vector j is 90°so vector j.vector k=j×k cos 90=-i and in vector k.vector k =k×k cos 0° =0